SOLVE AND CHECK THE ANSWER
A CONSULTANT TRAVELED IN A CAR FOR 3 HOURS TO ATTEND A MEETING. THE RETURN TRIP TOOK ONLY 2 HOURS (DIFFERENT ROUTE) BECAUSE HIS SPEED WAS 8 MILES PER HOUR FASTER. THE TOTAL TRIP WAS 291 MILES. WHAT WAS THE CONSULTANT'S SPEED EACH WAY?

Answer by ankor@dixie-net.com(12684)

A consultant traveled in a car for 3 hrs to attend a meeting. The return trip took only 2 hrs (different route) because his speed was 8 miles per hr faster. The total trip was 291 miles. What was the consultant's speed each way?
:
Let d = distance traveled for 3 hrs
and
(291-d) = distance traveled for 2 hrs (return trip)
;
Let s = speed which he traveled for 3 hrs
and
(s+8) = speed which he traveled for 2 hr (return trip)
:
Write two distance equations: Time * speed = distance
:
3s = d
2(s+8) = (291-d)
:
Substitute 2s for d in the above equation and find s:
2(s+8) = 291 - 3s
2s + 16 = 291 - 3s
3s + 2s = 291 - 16
5s = 275
s = 275/5
s = 55 mph on the way out
and
55 + 8 = 63 mph on the way back
:
:
Check solutions by finding total distance
Out distance = 3*55 = 165
Ret distance = 2*63 = 126
-------------------------
Total distance = 291